### v1

# Revised Mass and Orbit of Eridani b: A 1 Jupiter-Mass Planet on a Near-Circular Orbit

Abstract
The mature Jovian planet Eridani bĀ orbits one of the closest sun-like stars at a moderate separation of 3.5 AU, presenting one of the best opportunities to image a true analog to a solar system planet. We perform a thorough joint reanalysis and cross-validation of all available archival radial velocity and astrometry data, combining data from eight radial velocity instruments and four astrometric sources (Hipparcos, Hubble FGS, Gaia DR2, and Gaia DR3). We incorporate methodological advances that impact our findings including a principled treatment of correlation between Gaia DR2 and DR3 velocity and corrections for the changing light-travel time to this high proper motion system. We revise the planetās mass upward to and find that its orbit is nearly circular and close to coplanar with the outer debris disk. We further present one of the first models of an exoplanet orbit exclusively from absolute astrometry and independently confirm the planetās orbital period. We make specific predictions for the planetās location at key imaging epochs from past and future observing campaigns. We discuss and resolve tensions between previous works regarding the eccentricity, inclination, and mass. Our results further support that Eridani bĀ is one of the closest analogs to a Solar System planet yet detected around a nearby star.
ā±
1 Introduction
EridaniĀ is one of the closest planetary systems to Earth, just pc away. This proximity has made it a compelling target for direct imaging, giving us an opportunity to potentially image a true analogue to a planet in our Solar System: a relatively mature (200-800 Myr old, Mamajek & Hillenbrand (2008); 439 or 659 Myr old, Barnes (2007)) near-Jupiter mass planet, orbiting a Sun-like star at a separation that would fall between Mars and Jupiter ( AU).
This exosystem has been extensively studied since the 1980s, when pioneering Doppler radial velocity (RV) surveys by Campbell etĀ al. (1988) detected a long-term trend, and then Hatzes etĀ al. (2000) confirmed periodic motion of the star. Further credence was given to these RV signals when Benedict etĀ al. (2006) determined a complete orbit by combining RV data with astrometry from the Hubble Telescopeās Fine Guidance Sensor (FGS), the Allegheny Multichannel Astrometric Photometer (MAP) (Gatewood, 1987, 2000), and Hipparcos (van Leeuwen etĀ al., 1997), making it one of the earliest planets with an orbit derived in-part from absolute astrometry.
Since this time, studies of Eridani bĀ have entered a renaissance, driven by the promise that high-contrast direct imaging technology has advanced to the point where we may, in the near future, be able to study the planet directly via its thermally emitted light.
Several major attempts to image this system in the near-infrared have been published to date, including K-band imaging at Keck by Macintosh etĀ al. (2003), and M-band imaging with MMT by Heinze etĀ al. (2008), L-band imaging with VLT by Janson etĀ al. (2008), L- and M-band imaging with Spitzer by Janson, Markus etĀ al. (2015), M-band imaging with Keck by Mawet etĀ al. (2019), M-band imaging with Keck by Llop-Sayson etĀ al. (2021), and N-band at VLT imaging with Pathak etĀ al. (2021). Though these efforts did not achieve an imaging detection, they placed upper limits of on the brightness of the planet, and therefore limit its mass to for a 400 Myr system age.
Motivated by these attempts, upcoming and completed observations from JWST using NIRCam, MIRI, and NIRSpec are poised to reach significantly better sensitivity. This renews the need to make specific predictions of the planetās position in order to determine the best epochs to observe the planet, measure its mass and thus expected brightness, and confirm the location of any observed point sources. In addition, certain observations could benefit from knowing the location of the planet in advance to adjust the detector placement to mitigate saturation, plan dither patterns, and so on.
The most recent works to re-examine the orbit of Eridani bĀ include Reffert & Quirrenbach (2011), Giguere etĀ al. (2016), Llop-Sayson etĀ al. (2021), Roettenbacher etĀ al. (2022), FritzĀ Benedict (2022), and Feng etĀ al. (2023). These works examined the orbit of Eridani bĀ using a mix of data from various RV sources, Hipparcos (van Leeuwen, 2007), Gaia DR2 (Gaia Collaboration etĀ al., 2018), and Gaia DR3 (Gaia Collaboration etĀ al., 2023), but arrived at somewhat inconsistent conclusions with regard to the planetās orbital eccentricity, inclination, and dynamical mass.
The goal of this paper is to determine the orbit of Eridani bĀ using a careful reanalysis of all available archival data. We combine the superset of all observations considered in previous works, including radial velocity data from eight RV instruments (APF, HIRES, HARPS, Lick, CFHT, CHIRON, EXPRES, and CES) and four astrometric data sources (Hipparcos, Hubble FGS, Gaia DR2, and DR3). Throughout, we take special care to account for systematics and second-order effects that become non-negligible for fast-moving stars over long time-scales, including rigorous propagation of the systemās barycentric motion, perspective acceleration, changing parallax, and changing light travel-time to the receding system. We compare and contrast the methods and results of several previous analyses, finding where they agree and identifying which assumptions had led to disagreements.
In anticipation of a successful imaging detection, Eridani bĀ can also serve as a test run for future deep imaging campaigns with JWST, Roman, or HWO. Some of these campaigns will depend on radial velocity and astrometry to determine the location and brightness of planets ahead of time. Due to the long integrations required to detect sub-Jovian planets, it is incumbent on the community to demonstrate the ability to robustly determine planet positions and brightness ahead of targeted observations. We hope that a correct prediction of the location of Eridani bĀ can demonstrate this capability.
We now begin this paper by describing the sources of data we use, how that data is treated in our models, and other details of our modeling approach. We then present results, broken down by different subsets of data. Finally, we discuss how our results compare to previous efforts and the implications of our work.
2 Data and Methods
2.1 Radial Velocity Data
We gathered radial velocity data from many sources, starting with some of the earliest RV measurements made in the 1980s by CFHT (Campbell etĀ al., 1988). We used the ācorrectedā data presented in Table 1. We used data from Lick gathered from Fischer etĀ al. (2014), and split it into four groups as noted by the authors: Lick (1), before November 1994; Lick (2), dewar 13; Lick (3), dewar 6; Lick (4), dewar 8. We used data from CES (Zechmeister etĀ al., 2013), both from the LC and VLC, and treated these data separately. We further included data from HARPS (Trifonov etĀ al., 2020), split between before and the 2015 upgrade (āpreā and āpostā respectively). We used all HIRES and APF RVs presented by Mawet etĀ al. (2019) and by Llop-Sayson etĀ al. (2021). We further added RVs from CHIRON reported by Giguere etĀ al. (2016) and from EXPRES reported by Roettenbacher etĀ al. (2022).
This gives nearly complete coverage of the orbit for 40 years, between 1981 and 2021, in thirteen separate groups of RV data.
Over such a long time-scale, 3D perspective effects caused by the starās high proper motion ( 1ā/yr) and barycentric radial velocity (16 km/s) have the potential to build up to detectable levels. We elected to handle this apparent acceleration term directly in our models and thus attempted to undo any pre-existing static corrections applied to each dataset. We assumed that this correction was not applied for the CFHT, CES, HIRES, CHIRON, and EXPRES datasets and assumed that the correction was applied for the HIRES, APF, HARPS, and Lick datasets. For the latter datasets, we added back in secular acceleration at the rate given by Gaia DR3. To be conservative, we nonetheless include a linear trend for each dataset in our model. We encourage authors to include descriptions of these corrections in their works so that future authors can make the best use of their data, especially for systems with less orbital coverage than Eridani.
To model the RV data, we used a separate offset (), jitter (), and slope parameters () for each group, adding 39 variables to our model.
We modeled the activity signal from the star, which has been well established as having an signal from rotational modulation of the star and having an approximately month-long decay timescale.
We use the Celerite (Foreman-Mackey etĀ al., 2017) framework through the Celerite.jl
package111https://github.com/ericagol/celerite.jl/ and define an approximate quasi-periodic GP kernel with parameters , , , and , just as in Roettenbacher etĀ al. (2022).
We modeled each dataset with an independent Gaussian process, and multiplied their likelihoods. This treatment inherently assumes that interleaved data from different instruments are uncorrelated (see Blunt etĀ al. (2023), section 4). As discussed in Section 4.3, we tested the impacts of this assumption by comparing against a model with a single GP shared across all instruments, ultimately finding that our results in this case are not sensitive to the exact treatment.
2.2 Astrometric Data
The EridaniĀ system has been monitored for astrometric motion and parallax for over a century. We include precision absolute astrometry from Hipparcos (van Leeuwen, 2007), Hubble Fine Guidance Sensor (FGS, from Benedict etĀ al., 2006), and both Gaiaās second (DR2, Gaia Collaboration etĀ al., 2018) and third data releases (DR3, Gaia Collaboration etĀ al., 2023).
Beginning with Hipparcos, we used data from the mission in two different ways. We used the Hipparcos IAD epoch astrometry sourced from the Java Tool following the guidance of Nielsen etĀ al. (2020), renormalizing the uncertainties, and testing that we could reproduce the reported catalog measurements and uncertainties exactly. These data were used for the Hipparcos IAD model to be described shortly. We also used the Hipparcos proper motion sourced from the DR3 Hipparcos-Gaia catalog of accelerations (HGCA; Brandt, 2021), calculated from a mix of the 1997 and 2007 Hipparcos data releases, and calibrated against the Gaia DR3 velocity reference frame. These data were used in all other relevant astrometry models to ensure the reference frame was well-calibrated against that of Gaia DR3. This is a key departure from Feng etĀ al. (2023), who did not use the HGCA calibration between Gaia and Hipparcos reference frames, electing to include the reference frame alignment as variables in their model.
The second source of absolute astrometry data we used for this system was that of Benedict etĀ al. (2006). The data used in that publication were derived by measuring the positions of EridaniĀ using the Hubble Fine Guidance Sensor (FGS) relative to a field of nearby background stars. Although the raw data are available, the re-reduction of these data is out of scope for this work. Instead, we opted to extract the residuals from their parallax and proper motion fit (Figure 7 of Benedict etĀ al. (2006)) using Web Plot Digitizer (Rohatgi, 2025).
The extracted residuals are relative to a five-parameter linear model of annual parallax and proper motion fit to a combination of the FGS data, lower-precision astrometry from the Allegheny Multichannel Astrometric Photometer (not considered in this work). We justify using these residuals in our model by noting that the authorsā best-fit five-parameter model arrived at a parallax value that is very close to the measurement from Gaia DR3, meaning that no annual cyclic motion should be detectable in the residuals. Then, any residual proper motion not captured by their best-fit five-parameter model would appear as a linear trend. Such a linear trend could also be caused by frame rotation versus the Gaia DR3 reference frame, which we must assume is present to some degree. Rather than attempt to calibrate the absolute astrometry against Gaia with sufficient accuracy, we elect to simply fit linear trends to the residuals and fit only the acceleration over the FGS observations, which will be the same in any inertial reference frame. This is analogous to including a linear trend in models of RV data. This neatly sidesteps the issue of determining an absolute calibration of the FGS data and allowing us to use the residuals directly.
The final sources of absolute astrometry are both from Gaia. We include Gaia DR3 proper motion and uncertainties from the HGCA, just as we do the Hipparcos and Gaia-Hipparcos long-term proper motion. We undo the non-linear corrections incorporated by the HGCA to account for perspective acceleration and curvature in the R.A. and Dec. coordinate system, in favor of handling this motion inside our model.
We further include the Gaia DR2 catalog proper motion in some models. The DR2 and DR3 velocities can provide acceleration constraints because, despite sharing the majority of data between the two catalogs, the DR2 catalog values covers a time baseline that is shorter. Thus, despite the overlapping source data, a change in the velocity reported by the two catalogs can still provide some modest constraints on the acceleration of the star.
We model the two catalogs by simulating Gaia observations at each visibility window reported by GOST 222https://gaia.esac.esa.int/gost/index.jsp. We approximate the data using a single measurement per visibility window rather than per scan, just as in most previous works. We then perform a linear least-squares fit of a 4-parameter model to both the DR2 time-range and the DR3 time-range. We then compared the velocities extracted from these simulations, against the reported catalog values .
This is similar in principle to the work of (Feng etĀ al., 2023), though we differ in how the data are modeled a few important respects. First, we handle the correlation between DR2 and DR3, and second, we use the HGCAās spatially dependent calibration between the Hipparcos and Gaia reference frames and calibrated (inflated) uncertainties rather than those reported by Gaia directly.
The reason we must model the correlation between DR2 and DR3 is that the reported velocities are derived from as much as of the same measurements. We therefore form a full covariance matrix between R.A. and Dec. velocities from both catalogs as the block matrix:
While the DR2 and DR3 catalogues report the covariance between derived R.A. and Dec. velocities, and respectively, we must determine an appropriate form for the inter-cataloge covariance .
We estimate the off-diagonal blocks of the covariance matrix using the reported uncertainties from Gaia DR2 and DR3 and the fraction of shared epochs. We assume that the sampling is uniform in time. Under these assumptions, consider a 1-D case with uniform time sampling, and the calculation of slope of a set of independent Gaussian random variables with uniform uncertainties, and the slope of a set of N independent Gaussian random variables which include the same set of (where ) of random variables as in (just as Gaia DR3 includes all measurements from Gaia DR2, plus an addition set of new measurements).
To determine the correlation between and , we start with the least squares slope of a line which is given by
This can be rearranged to show that the slope is a linear combination of the y-values:
where
Therefore, our two slopes can be written as:
and
where are independent normal random variables with variance . The variances are:
and
and the covariance is
Giving, therefore, the corelation
This demonstrates that the correlation between the two slopes depends only on the proportion of shared variables, under these idealized assumptions. We therefore treat the Gaia DR2 proper motion in either R.A. or Dec. as having the correlation with the proper motions reported by DR3.
The last detail to complete is the cross terms giving covariance between the DR2 proper motion in R.A. and DR3 proper motion in Dec., and vice-versa.
We simply generalize the covariance matrix between two scalar variables,
to the matrix case
where as above and where and are the matrix square roots, which are unique for positive-definite matrices like covariance matrices.
One additional consideration is that the DR2 proper motion reference frame is known to be rotating slightly compared to DR3. We corrected for this according to from Lindegren, Lennart (2020). Unlike previous works that simply adopted the mean values of these parameters, we included them as priors in our models in order to rigorously marginalize over the uncertainties in the reference frame calibration. This is similar in effect to the offsets employed by Feng etĀ al. (2023), but only applied to the DR2-DR3 velocity difference.
As noted by the Gaia teams and in the HGCA (Gaia Collaboration etĀ al., 2016; Brandt, 2018, 2021), the formal uncertainties reported by Gaia and Hipparcos are underestimated. Accordingly, the the HGCA catalogs inflate the proper motion uncertainties such that a sample stars that are not likely to be accelerating have standardized residuals that follow a Gaussian distribution. This error inflation factor for Gaia DR2 and DR3 varies with spatial position. Where Feng etĀ al. (2023) adopts a flexible jitter parameter, we used the calibrated uncertainties from the HGCA catalogs for both DR2 and DR3.
Based on all these considerations, the covariance matrix we adopt between Gaia DR2 and DR3 R.A. and Dec. proper motions is
which corresponds to the correlation matrix
We note the high correlations, as expected, between DR2 and DR3 proper motions along the same axes, moderate correlation between the proper motions along different axes in the same data release, and low correlation between the axes of the different data releases.
2.3 Imaging
There is raw image data available from previous campaigns, including those by Mawet etĀ al. (2019), Llop-Sayson etĀ al. (2021), and Pathak etĀ al. (2021). We did not include these images in our model since the published sensitivity curves indicate that the planet would be below the level of the noise for reasonable inclination angles (limit of from Mawet etĀ al. (2019), 2ā4 Ā from Pathak etĀ al. (2021)). A future work could explore integrating image data into this orbit model following the same recipes as Llop-Sayson etĀ al. (2021) and Thompson etĀ al. (2022), especially if more images of comparable or greater sensitivity become available.
2.4 Non-Linear Effects
In order to rigorously account for second order acceleration terms originating from the systemās close proximity to Earth, we model the three-dimensional motion of the systemās barycenter through space.
We convert the the proper motion, parallax, and radial velocity at a reference epoch, taken to be J2016, into Cartesian coordinates and propagate them linearly versus time assuming uniform space motion. This accounts for perspective (secular) acceleration, for curvature in the R.A. and Dec. spherical coordinate system, and for changing parallax versus time which causes an slight apparent magnification versus time effect.
In addition, we identify that that changing light travel-time is an important and previously neglected consideration for this systemāand perhaps many others. RV and pulsar timing data are often corrected for the so-called āRĆømer-delay,ā the changing light travel time between a source and the Earth due to the Earthās travel around the Sun. In addition to the Earthās motion, we note that the distance from our solar system to EridaniĀ has increased by 19 light-hours between 1980 and 2020. This has a small effect on the apparent orbital motion of the planet on the order of 1 mas, and on the perturbed stellar radial velocity on the order of 2 cm/s, but it has a significant effect on the observed long-term proper motion: 975 mas/yr 19 hr = 2.1 mas less displacement than expected from straight-line motion extrapolated over this time span. This is larger than the astrometric uncertainties of missions like Gaia and Hipparcos.
In the context of the Hipparcos-Gaia position difference, the apparent time delay of 11.5 hr corresponds to an apparent long term proper motion that is 0.05 mas/yr lower than a naive model would suggest, an effect greater than the uncertainty reported by the HGCA. Precision orbit models of high proper motion stars should take this effect into account when modeling long-term proper motion.
To account for this changing light travel time in our models, we solve for the emission time at each epoch that corresponds to our observation time using an iterative scheme. We calculate this difference compared to the reference epoch J2016, for which we set ; that is, we donāt consider the total light travel time, only the change in light travel time from the systemās distance at the reference epoch.
Lastly, we do not consider the third order perturbations to this motion caused by the planet; that is, the above effects are calculated for the systemās barycenter applied to both the star and planet.
We make these corrections at every epoch and model evaluation since the barycentric proper motion of the system is allowed to vary, and projecting the (near-) instantaneous proper motion from Gaia back would unintionally mix in contributions from the planetās orbit.
2.5 Gaussian Process
EridaniĀ has significant photometric and RV variability consistent with a rotation period of 10-12 days. As with the works of Llop-Sayson etĀ al. (2021), Giguere etĀ al. (2016), and Roettenbacher etĀ al. (2022), we model this activity signal using a Gaussian process with a quasi-periodic kernel. We implemented the GP in our model using the Celerite.jl package from Foreman-Mackey etĀ al. (2017), after forking and updating the package to work in modern Julia. We fit the GP separately to each dataset, but share a single set of GP hyper-parameters between all instruments.
6em
2.6 Implementation
We implemented our models in the Julia-based (Bezanson etĀ al., 2012) Octofitter framework (Thompson etĀ al., 2023) and sample using stabilized variational non-reversible parallel tempering Surjanovic etĀ al. (2022); Syed etĀ al. (2021) implemented in Pigeons.jl (Surjanovic etĀ al., 2023).
Unlike the model presented in Llop-Sayson etĀ al. (2021), and like the models presented in Feng etĀ al. (2023) and FritzĀ Benedict (2022) we jointly model both the astrometry and radial velocity, rather than model the RV and astrometry separately. Unlike Feng etĀ al. (2023) and FritzĀ Benedict (2022), we include the Gaussian process model of stellar activity. As with FritzĀ Benedict (2022) we include the Hubble FGS astrometry originally published in Benedict etĀ al. (2006).
The largest model which incorporates every listed data source has a total of 61 different parameters. The majority of these are instrument-specific nuisance parameters, which we marginalize over when presenting results. The full list of parameters and our adopted priors are listed in Table 1. Whenever we sample from a model that ignores a certain dataset, we remove the related instrument-specific nuisance parameters (e.g. a model without any RV data does not include the many RV offsets, jitters, and slope parameters).
We sample each model until the log-evidence converges between sampling rounds and the diagnostic (Gelman & Rubin, 1992; Vehtari etĀ al., 2021) is to at least two or more decimal places.
Parameter | Description | Unit | Prior |
Primary mass | |||
Secondary mass | Uniform(0, 10) | ||
Orbital period | yr | Uniform(5.36, 9.36) | |
Orbital eccentricity | Uniform(0, 0.999) | ||
Orbital inclination | deg | Sine | |
Longitude of ascending node | deg | Uniform(0,360) | |
Orbit fraction at reference epoch | Uniform(0,1) | ||
Argument of periastron | deg | Uniform(0,360) | |
GP amplitude | m/s | Uniform(0.00001, 2000000) | |
GP amplitude ratio | Uniform(0.00001, 200) | ||
GP decay timescale | days | Uniform(2, 200) | |
GP rotation period | days | Uniform(8.5, 20) | |
APF zero point | m/s | ||
EXPRES zero point | m/s | ||
CHIRON zero point | m/s | ||
CFHT zero point | m/s | ||
HIRES zero point | m/s | ||
Lick (1) zero point | m/s | ||
Lick (2) zero point | m/s | ||
Lick (3) zero point | m/s | ||
Lick (4) zero point | m/s | ||
HARPS (pre) zero point | m/s | ||
HARPS (post) zero point | m/s | ||
CES (lc) zero point | m/s | ||
CES (vlc) zero point | m/s | ||
APF jitter | m/s | Uniform(0.1, 100) | |
EXPRES jitter | m/s | Uniform(0.1, 100) | |
CHIRON jitter | m/s | Uniform(0.1, 100) | |
CFHT jitter | m/s | Uniform(0.1, 100) | |
HIRES jitter | m/s | Uniform(0.1, 100) | |
Lick (1) jitter | m/s | Uniform(0.1, 100) | |
Lick (2) jitter | m/s | Uniform(0.1, 100) | |
Lick (3) jitter | m/s | Uniform(0.1, 100) | |
Lick (4) jitter | m/s | Uniform(0.1, 100) | |
HARPS jitter | m/s | Uniform(0.1, 100) | |
CES (lc) jitter | m/s | Uniform(0.1, 100) | |
CES (vlc) jitter | m/s | Uniform(0.1, 100) | |
APF slope | m/s/day | ||
EXPRES slope | m/s/day | ||
CHIRON slope | m/s/day | ||
CFHT slope | m/s/day | ||
HIRES slope | m/s/day | ||
Lick (1) slope | m/s/day | ||
Lick (2) slope | m/s/day | ||
Lick (3) slope | m/s/day | ||
Lick (4) slope | m/s/day | ||
HARPS (pre) slope | m/s/day | ||
HARPS (post) slope | m/s/day | ||
CES (lc) slope | m/s/day | ||
CES (vlc) slope | m/s/day | ||
DR2 velocity offset, x | mas/yr | Normal(-0.068, 0.051) | |
DR2 velocity offset, y | mas/yr | Normal(-0.051, 0.042) | |
DR2 velocity offset, z | mas/yr | Normal(-0.014, 0.036) | |
Parallax | mas | Uniform(310.577, 0.135) | |
Proper motion in RA | mas/yr | ||
Proper motion in Dec | mas/yr | ||
Fine Guidance Sensor jitter | mas | LogUniform(0.1, 100) | |
FGS offset in | mas | Uniform(-500,500) | |
FGS offset in | mas | Uniform(-500,500) | |
FGS trend in | mas/d 7 | Uniform(-10,10) | |
FGS trend in | mas/d 7 | Uniform(-10,10) |
12em
2.7 List of Models
In order to assess the agreement between different datasets and rule out systematic errors, we create a set of 7 different models. We progressively add additional data to build up to the complete model. We compare the parameters inferred by each model and their predictions for the planetās location to study what constraints come from which dataset. Checking that the inferred parameters and predictions are not in tension between models will allow us to check for signs of systematic errors that could bias our results.
The set of seven main models we consider are:
1.
All RV data (no astrometry)
2.
All astrometry data (no RV)
3.
RV and Hubble FGS astrometry
4.
RV and Hipparcos IAD (capturing just the motion within the Hipparcos mission)
5.
RV and Hipparcos-Gaia acceleration from the HGCA (Brandt, 2021) (capturing the net motion at the Hipparcos and Gaia epochs, versus their long-term trend between missions)
6.
RV and HGCA, with the addition of the Hubble FGS astrometry
7.
Full model with RV, HGCA, FGS, and the DR vs. DR3 acceleration.
The 1st (RV only) and 2nd (astrometry only) models include entirely separate data, aside from the position and parallax of the star from Gaia DR3. Similarly, if we consider that the orbital semi-major axis, eccentricity, and argument of periastron are determined largely by RV data, we can then also compare the nearly independent constraints on the inclination and position angle of ascending node from the Hipparcos IAD, the Gaia-Hipparcos acceleration, and the Hubble FGS data using results from the 3rd, 4th, and 5th models.
-1.45cm-1.2cm
-1.8cm-0.3cm
3 Results
We now present the results of our models, starting with only the RV data and building up to the final solution. The posteriors from all seven models are compared in Figure 1 and 2, which includes a corner plot and predictions for 2025-01-01. We use consistent colors to identify each model throughout the figures in this section. We report results using a mix of credible intervals (CI) and highest density intervals (HDI) as appropriate to summarize peaked or truncated distributions, with the notation used to indicate the median and 14% / 84% quantiles, and the [lower,upper] notation to indicate the the 75% HDI interval. As with other works, we use 75% to emphasize the arbitrariness of the specific probability mass fraction. Tables summarizing the posteriors of each model are presented in the Appendix.
3.1 RV Data
Our model of the RV data alone results in a posterior that is unimodal, with a well-constrained semi-major axis of AU and a semi-amplitude of m/s, which is consistent with previous works. The maximum a-posteriori orbit sample from the RV-only posterior is displayed in Figure 3. The orbital phase is well determined. The RV data and parallax alone show that the planet will be at opposition (its maximum separation from the star) in 2017, 2021, 2025, 2029, and so on. During these times, the planet will be localized to a ring of mas separation. The minimum mass, , is . The RV data supports an eccentricity that is consistent with 0: the 75% highest density interval is .
The GP hyper-parameters adopt values similar to previous works, with a rotation period, , of d, and decay timescale, , of d.
3.2 Constraints added by the Hipparcos IAD
Adding the Hipparcos IAD data provides constraints on the inclination, position angle of ascending node, and position angle of the planet versus time. This constrains the planet to be roughly on the South side of the star in 2025, North in 2029, and so on. It provides essentially no information on the horizontal position of the planet. The addition of the Hipparcos IAD does not significantly constrain the mass of the planet, other than that it is less than 1.3 Ā (84 percentile upper limit), or the 75% HDI of [0.56, 1.09] .
In the versus subpanel of Figure 2, we see that the Hipparcos IAD data does not provide any strong constraints on the inclination of the planetās orbit, but it does suggest that if the planet is co-planar with the outer disk, then it is more likely to orbiting counterclockwise.
3.3 Constraints added by the HGCA
Adding the DR3 Hipparcos-Gaia proper motion anomaly from the eDR3 HGCA (Brandt, 2021) to the RV data provides significantly stronger constraints. Note that this model includes only net velocities from the Gaia and Hipparcos cataloges, and does not include the within-catalogue motion from the Hipparcos IAD. With RV and the HGCA data, we find a mass upper limit (84th percentile) of 1.00 . With regard to the orientation of the orbit, the HGCA data constrains it to a bimodal distribution of either or , and a position angle of ascending node of , both of which are in agreement with the orientation of the outer debris disk to (Booth etĀ al., 2023).
From the RV and HGCA data, we can constrain the position angle of the planet to be SSW of the star in 2025 in a region only about 600mas wide.
Figure 4 shows the RV, proper motion, and position perturbation in R.A. and Dec. versus time for different orbit draws. Looking at the proper motion panels, the orbit is near its most extreme values in the Dec. direction, but at mid-phase in R.A.āhence why the motion in Dec. is well-determined but ambiguous in R.A. with these data alone.
3.4 Constraints added by the Hubble FGS data
Adding the Hubble FGS data from Benedict etĀ al. (2006) independently constrains the planet to be SSE or S of the star in 2025. These data prefer a slightly lower 84% upper mass limit of 0.89 .
Looking at Figure 4, we see that the addition of the FGS data to the RV and HGCA removes an ambiguity in R.A. motion of the star, while agreeing well with the Dec. motion already determined by the HGCA.
We see that the curvature in the FGS data is in very good agreement with the sinusoidal model, though recall that the absolute value and linear trend are fit flexibly as part of the model.
3.5 Constraints added by the Gaia DR3-DR2 acceleration
Similarly to the addition of the FGS data, adding the DR3-DR2 acceleration removes an ambiguity in the R.A. motion of the star. Notably, the constraint provided by either the DR3-DR2 acceleration or the Hubble FGS astrometry selects the same part of the ambiguous posterior from the RV & HGCA alone, increasing our confidence in the model.
3.6 Complete Model
With the complete model, the planet is very well-localized (Figure 2, top right panel). The model yields a semi-major axis of AU, a mass of , eccentricity of [0.00,0.09], and an orbit that is inclined . The direction of motion is counterclockwise in the plane of the sky.
We find that all proper motion data are well-fit by the model, as shown in Figure 5.
3.7 Astrometry only orbit solution
The data from Hipparcos, Hubble FGS, Gaia DR2, and DR3 is sufficient to, for the first time, fit an astrometry-only model, without any RV data. Using a uniform prior on the orbital period of 1ā10 yr, we find an orbit posterior with a semi-major axis of AU. We find the model prefers moderately low eccentricity solutions (HDI of [0, 0.42], peaked at 0). The orbital period agrees to better than with the RV only model. The times of conjunction and opposition are likewise moderately well constrained, though not as precisely as by the RV data. The astrometry data independently constrains the mass of the planet to .
The posterior is bi-modal in position angle of ascending node, though both modes could agree with the orientation of the outer debris disk to (the orientation of a disk is ambiguous to in inclination and in position angle of ascending node with the current data).
3em
4 Discussion
4.1 Orbital Period
All of our orbit models that include RV data support an orbital period in the range of 7.3ā7.45 years, with the complete model giving yr. This is very similar to the period of yr reported by Llop-Sayson etĀ al. (2021), though somewhat shorter than the yr or yr periods reported by Feng etĀ al. (2023) and by Benedict etĀ al. (2006) respectively. We believe this disagreement stems from differences in the noise model used to deal with stellar activity, which we discuss in Section 4.3.
Our model of the absolute astrometry only, with no RV data, supports a period of years. This independent support for the orbital period should lay to rest concerns that the RV signature of Eridani bĀ may be an artifact of long-term stellar activity cycles.
Indeed, the perturbation orbit of the star has a projected semi-major axis of 2.7 mas which is just over twice the 1.1 mas apparent radius of the star (using a 0.738 stellar radius from Rains etĀ al. (2020)). Such a large displacement of the photocenter over a period of years cannot be explained by stellar activity.
4.2 Orbital Phase
The RV model we present determines the overall phase of the orbit and constrains the times of conjunction and opposition to approximately one month. Given the known parallax of the system, this means that the RV data alone can constrain the separation of the planet to ring whenever it is near oppositionāin 2017, 2021, 2025, 2029, and so on. This was noted as well in Mawet etĀ al. (2019).
The work by Llop-Sayson etĀ al. (2021) added constraints from Hipparcos and Gaia DR2 and presented predictions for the planetās location at two imaging epochs. The astrometry data added constraints to the PA versus time, but the additional data broadened the constraints to the planetās separation compared to RV data alone. We hypothesize this may be a result of the two-step process in which the RV data were modeled first, a kernel density estimate (KDE) was fit to the posterior, and then the KDE was used as a prior in the subsequent absolute astrometry fit. The variables used in the astrometry fit do not include the time of conjunction directly, expressing it instead as a combination of other variables. It could be that while the KDE accurately represented the marginal posteriors of the period, eccentricity, argument of periastron, and epoch of periastron passage, the detailed correlation structure may not have been captured with sufficient detail to give a tightly constrained prior on time of conjunction derived from a combination of the above.
Joint models like the ones presented here and in previous works Benedict etĀ al. (2006); Reffert & Quirrenbach (2011); FritzĀ Benedict (2022); Feng etĀ al. (2023) sidestep this challenge. This could explain why our RV and Hipparcos model has considerably less uncertainty on the separation of the planet near the times of conjunction (e.g. in 2017-09-01, see Figure 9).
6em
4.3 Eccentricity
The eccentricity of the planet is a key concern since it has a strong impact its separation from the star at opposition. Our models all support a low eccentricity that is consistent with zero. This is consistent with previous works that included a GP model of stellar activity (Mawet etĀ al., 2019; Llop-Sayson etĀ al., 2021; Giguere etĀ al., 2016) (eccentricity near zero) and is inconsistent with works that did not such as Feng etĀ al. (2023) and FritzĀ Benedict (2022) (eccentricity of 0.26 and 0.16 respectively).
6em
4.3.1 Including a GP results in a low eccentricity
To determine the source of this disagreement, we compared an RV model with a GP and an RV model without a GP. As expected from earlier studies, the model without a GP arrives at an eccentricity of , while the model with a GP supports a low eccentricity consistent with zero.
6em
4.3.2 The GP results are robust to the choice of data
-1.0cm-0.3cm
To understand why these two models arrive at different results, we performed two suites of leave-one-instrument-out cross validation. For each model, we created 13 copies that each drop one of the instrument datasets (where e.g. Lick (1) and Lick (2) are treated as separate datasets). For this test only, we bounded the eccentricity to be at most 0.55. We then sampled the posterior for all 26 models and compared their marginal eccentricity posteriors, shown in Figure 6.
We find that the GP model produces eccentricity posteriors that are robust to the choice of dataset dropped. In all cases, we arrive at a probability distribution that is consistent with a near-circular orbit. The removal of either the APF or HIRES data broadens the eccentricity posterior but does not pull it significantly away from zero. This indicates that the APF and HIRES data provide the strongest upper bounds on the eccentricity, but does not indicate any tension from these datasets.
The no-GP model, on the other hand, is sensitive to the choice of dataset dropped. There are several datasets (HARPS (pre), Lick (4), and APF) that, when held out, lead to a drastic shift in the reported eccentricity. Most notably, dropping the HARPS (pre) data results in an eccentricity posterior with a 75% HDI of [0.035, 0.138] ( CI), compatible with zero.
While our main models used separate GPs for each instrument, we also tested a model with a single GP for all data. This setup, related to a strategy suggested in Blunt etĀ al. (2023), shares temporal correlation information between instruments, instead of treating data from each instrument as uncorrelated. We find essentially the same low-eccentricity results from this model, indicating that the GPs do not appear to be overfitting the data from each instrument.
6em
4.3.3 High cadence data are the source of the disagreement
-1.25cm-0.5cm
The HARPS (pre) series contains a burst of high cadence data, with 430 out of all 530 points captured over a 2-night range in October, 2004. This dense data from HARPS visibly traces out a correlated signal with a timescale much higher than the Keplerian signal (Figure 7), which we interpret as a star-spot moving into or out of view following the 11.68 day rotation period of the star (Barnes, 2007). We find that dropping this data from the no-GP model leads to an eccentricity distribution that essentially matches that of our GP model. The no-GP model is not equipped to recognize the high correlation between these points, and so the burst of high cadence data distorts the orbit modelāproducing both a non-zero eccentricity, and an overconfident posterior.
6em
4.3.4 Possible impacts to previous RV models
Feng etĀ al. (2023) used a moving average (MA) model to account for temporal correlation in the noise, which can be thought of as an simplified GP.
The MA model described in that work uses the past data points to predict the subset point, with for the HARPS (pre) data. We hypothesize that the MA model may be affected in a similar way to the no-GP model when it encounters a burst of hundreds of data points over a short period of time. Earlier data points containing different realizations of the stellar noise may fall outside of this -point memory, which could limit the models ability to recognize the temporal correlation and result in a higher eccentricity. It is possible that the MA model might arrive a low eccentricity when these points are dropped, just like our no-GP model.
4.3.5 Comparison with astrometry-only models
It was noted by Feng etĀ al. (2023) that ā..astrometry-only analyses have also resulted in high-eccentricity solutionsā¦.ā While the cited works of Benedict etĀ al. (2006) and FritzĀ Benedict (2022) did arrive at moderate eccentricities, we should clarify that they did so using joint models of RV and astrometry which likewise did not include a GP or other model of the stellar activity. In fact, no previous works have published an astrometry-only analysis to the best of our knowledge. In contrast, our true astrometry-only model finds an eccentricity with a 75% highest density interval of [0, 0.44], peaking at zero. The astrometry data therefore can be compatible with either the low or moderate eccentricity results.
4.3.6 Conclusion: support for low eccentricity
We find that the GP models do not appear to be over-fitting a real eccentricity signal, and in fact are more predictive of held out data than the simpler model not including a GP. Based on the improved predictive performance of the GP model, we conclude that the current data supports a low eccentricity orbit, consistent with zero to the precision of currently available data, in agreement with Mawet etĀ al. (2019), Llop-Sayson etĀ al. (2021), and Roettenbacher etĀ al. (2022).
4.4 Direction of Orbital Rotation
We find that our models of individual astrometric data sources do not strongly constrain the direction of apparent orbital rotation in the plane of the sky to be clockwise or counterclockwise. For instance, the model of the Hipparcos and Gaia DR3 proper motions is evenly split between the two directions (inclination below/above ). The model of the Hipparcos IAD does suggest that either the planet is moving counterclockwise and is near-coplanar with the disk, or, that the planet is moving clockwise but misaligned with the disk. That said, the combination of all data sources, particularly the Hubble FGS and Gaia DR3-DR2 acceleration, supports counterclockwise motion fairly conclusively (see Figure 4). This result is inconsistent with that of Feng etĀ al. (2023) which is clockwise as indicated in their Figure 2.
6em
4.5 Coplanarity with outer debris disk
We now compare our complete orbit model to the outer debris disk model fit to ALMA data by Booth etĀ al. (2023). We find that the HGCA, Hubble FGS, and DR3-DR2 all support a planet that is co-planar with the outer debris disk to in both inclination and position angle of ascending node. Investigating our models, we find that the data point pulling the model slightly away from co-planar is the Gaia-Hipparcos position difference. We hypothesize that if the planet is indeed co-planar, then a very slight miscalibration between the Hipparcos and Gaia velocity reference frames (as corrected by the HGCA), or a slight inaccuracy in our model of the HGCA position difference would account for this slight tension. This data point is particularly sensitive to systematics because of the very high proper motion of the star surfacing non-linearities in the R.A.-Dec. coordinate system and variable light travel time to the system.
An alternative explanation might come from the disk model itself, which assumes a perfectly symmetric disk. If this assumption were violated by a sufficiently clumpy substructure, it is possible that the revised position angle of ascending node could shift a small amount (M. Booth, private communications).
Regardless, when we compare the inclination of the complete model with that of the outer disk from Booth etĀ al. (2023), we find the a mutual inclination 75% HDI of . This is nearly consistent with coplanarity but we cannot rule out a modest misalignment.
-1.2cm-0.45cm
4.6 Mass
The RV data strongly constraints the of the planet, such that the mass we report is essentially determined only by the inclination of the orbit. This situation is presented in Figure 8. We find that the full model prefers an inclination of , which corresponds to a mass of . If the planet is indeed perfectly coplanar with the outer disk to within , this value would rise slightly to .
This mass is substantially higher than the value of reported by Llop-Sayson etĀ al. (2021), but this disagreement is only due to the different inclination ranges we findāadjusting the inclination of their posterior down to would give a mass of as well.
The mass we find is also higher than the value of reported by Feng etĀ al. (2023), but in this case the disagreement comes partially from the different reported inclination () and partially from the reported non-zero eccentricity of 0.26.
Based on the revised mass, we calculated an expected contrast ratio of the planet versus star to be in L band, and in M band using Sonora Bobcat models (Marley etĀ al., 2021) and an age of 439 Myr (Barnes, 2007). That said, these are nominal values only as there is considerable uncertainty in the age of the system.
4.7 Predictions
We present predictions in Figure 9 of the planetās apparent location for every year from 2018 through to 2029, in addition to four significant imaging epochs. From the RV data alone, the planetās location is confined to a narrow ring of mas at opposition, which occurs in 2024-12-07 + yr. We the addition of any individual absolute astrometry dataset, the position angle of the planet at conjunction is confined to be South of the star in 2025, North in 2029, and so on. With our full model incorporating all absolute astrometry data sources, the position is confined to a region just mas wide near opposition. With the revised inclination, an imaging detection of the planet away from opposition may also be feasible, as we find that it never appears closer to the star than mas. This last constraint does rely on the complete model of all absolute astrometry, so conservative imaging programs may still wish to time their observations near opposition.
4.8 Expected Constraints from Relative Astrometry
With regards to the orbit and mass determination, the main benefit of an imaging detection will be an accurate measurement of the relative astrometry between the planet and star. An imaging detection near opposition, while the most technically feasible would not contribute significant constraints to the inclination and therefore mass but would serve to confirm the orbitās position angle of ascending node. To constrain inclination and mass, we recommend an imaging epoch somewhat away from opposition. The strongest constraints on the planetās mass would come from a relative astrometry measurement recorded close to a quarter period away from conjunction, i.e. when the planet is closest to the star. Such an endeavor would of course have to be balanced with the detectability at this closer separation.
6em
5 Conclusions
In this work we have presented an extensive reanalysis and cross-validation of all available radial velocity and absolute astrometry data. We determine the period of the planetās orbit to be yr ( AU). We independently confirm the period using the first complete orbit model of the system from only absolute astrometry data. In doing so, we rule out the possibility that the 7.3 year periodic signal could be a result of stellar activity cycles instead of a planet.
We find a much lower overall inclination than previous works (, ) that is likely close to coplanar with the outer debris disk (Booth etĀ al., 2023), with the 75% HDI of the mutual inclination between . The higher inclination we find supports a dynamical mass of . We also determine that the planetās orbit has a low eccentricity of consistent with a near-circular orbit.
We discussed disagreements in the literature regarding the planetās time of conjunction / opposition, inclination, and eccentricity, and presented arguments for why our updated results differ.
Finally, we presented detailed predictions for the planetās apparent location versus time, broken down by different subsets of data.
Our significantly revised mass and position predictions will aid future efforts to detect the planet through imaging, and point to Eridani bĀ being a remarkably close analog to our own planet Jupiter.
We thank T. Brandt for guidance in modeling the HGCA. We thank William Robertson for discussions regarding second order effects for high proper motion and barycentric radial velocity stars. We also thank Eric Agol for discussions regarding Celerite.jl and Gaussian processes. We thank Jessica Speedie for discussions regarding ALMA data. We thank Mark Booth for sharing posteriors from Booth etĀ al. (2023) in machine-readable format.
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
Appendix A Additional Tables
We now present summary statistics for each variable of our models. Table 2 presents the 75% HDI, while Table 3 presents the 68% CI.
The residuals we digitized from Figure 7 of Benedict etĀ al. (2006) are also presented in Tables 4 and 5. Based on that figure, we interpret the coordinate as a negative change in declination, while is a negative change in right ascension, times the cosine of declination.
Parameter | RV | HIP, FGS, DR2, DR3 | RV, FGS | RV, Hip. IAD | RV, HGCA | RV, HIP, FGS, DR3 | RV, HIP, FGS, DR2, DR3 |
Parameter | RV | HIP, FGS, DR2, DR3 | RV, FGS | RV, Hip. IAD | RV, HGCA | RV, HIP, FGS, DR3 | RV, HIP, FGS, DR2, DR3 |
epoch [MJD] | Position Residual [mas] |
51946.5 | -1.84392 |
51946.5 | -1.81306 |
51946.8 | -1.31929 |
51947.0 | -1.01754 |
51980.5 | -1.72745 |
51980.5 | -1.69316 |
51981.2 | -2.54699 |
51981.3 | -2.46126 |
51981.3 | -2.39268 |
51981.4 | -2.31724 |
51981.4 | -2.25209 |
51981.7 | -1.84747 |
51982.0 | -1.27826 |
51982.0 | -1.25769 |
51982.1 | -1.23369 |
51982.1 | -1.1171 |
51982.2 | -1.00051 |
51982.2 | -0.973082 |
51982.3 | -0.856497 |
51983.2 | -1.45314 |
52179.3 | -2.25963 |
52179.3 | -2.22819 |
52179.3 | -2.17962 |
52179.4 | -2.10532 |
52179.4 | -2.08532 |
52179.7 | -1.54811 |
52179.9 | -1.20236 |
52179.9 | -2.84827 |
52180.0 | -2.81112 |
52180.0 | -2.77683 |
52180.0 | -2.40365 |
52180.0 | -2.72826 |
52180.5 | -1.9996 |
52180.5 | -1.95674 |
52180.5 | -1.87387 |
52180.6 | -1.80815 |
52180.6 | -1.77386 |
52180.8 | -1.43096 |
52181.1 | -1.01091 |
52181.8 | -1.48525 |
52308.3 | -3.63452 |
52308.4 | -3.57451 |
52308.5 | -3.29734 |
52308.6 | -3.16018 |
52308.6 | -3.1316 |
52308.7 | -2.98016 |
52308.8 | -2.92872 |
52308.8 | -2.85443 |
52309.0 | -2.62297 |
52309.1 | -2.48581 |
52309.1 | -2.43152 |
52309.1 | -2.41723 |
52309.2 | -2.18006 |
52309.3 | -2.11148 |
52309.3 | -2.07148 |
52309.7 | -1.4714 |
52309.8 | -2.90586 |
52309.9 | -1.12279 |
52310.0 | -0.977059 |
52310.1 | -0.854187 |
52310.3 | -0.559866 |
52311.3 | -0.51415 |
52497.0 | -3.48657 |
52497.3 | -3.02366 |
52498.0 | -3.63802 |
52498.3 | -3.20654 |
52498.4 | -3.01224 |
52498.5 | -2.78364 |
52499.5 | -2.92937 |
52499.8 | -2.37502 |
52503.6 | -3.07798 |
52503.7 | -2.86081 |
52503.7 | -2.82937 |
52504.0 | -2.45504 |
52505.0 | -2.44647 |
52505.4 | -3.54375 |
52506.3 | -2.06929 |
52508.5 | -3.62092 |
52509.5 | -1.955 |
52510.6 | -1.88357 |
52512.9 | -3.27803 |
52513.8 | -3.56664 |
52514.1 | -2.978 |
52514.3 | -2.64653 |
52535.5 | -3.98962 |
52535.7 | -3.68959 |
52535.7 | -3.65816 |
52536.2 | -2.90092 |
52537.5 | -2.4323 |
52539.7 | -4.01821 |
52539.8 | -3.82962 |
52540.2 | -3.16097 |
52540.7 | -2.3523 |
52542.4 | -3.06096 |
52543.7 | -2.58091 |
52544.0 | -2.20086 |
52642.1 | -2.79556 |
52642.9 | -3.17846 |
52642.9 | -3.1156 |
52643.8 | -1.664 |
52645.5 | -2.28979 |
52646.2 | -2.97845 |
52646.4 | -2.51268 |
52646.5 | -2.43553 |
52646.6 | -2.23836 |
52649.4 | -2.8413 |
52650.1 | -3.40137 |
52650.5 | -2.72129 |
52651.5 | -2.7613 |
52673.0 | -3.68148 |
52673.6 | -2.71851 |
52673.6 | -1.03259 |
52673.7 | -2.57278 |
52674.2 | -1.65267 |
52678.4 | -3.3929 |
52678.8 | -2.6871 |
52679.1 | -2.30419 |
52679.4 | -3.42719 |
52680.8 | -2.91285 |
52682.2 | -3.95012 |
52682.3 | -3.81868 |
52683.1 | -2.52423 |
52683.3 | -2.21277 |
52683.8 | -1.41553 |
52684.0 | -1.12121 |
52685.9 | -3.13575 |
52686.5 | -3.82726 |
52686.7 | -1.86702 |
52686.8 | -3.33006 |
52688.2 | -2.78428 |
52689.7 | -1.98133 |
52689.8 | -1.87846 |
52860.9 | -3.11348 |
52862.1 | -2.87346 |
52864.3 | -2.65916 |
52864.6 | -3.80215 |
52865.6 | -2.35055 |
52866.5 | -2.49057 |
52869.4 | -2.92206 |
52869.6 | -2.59917 |
52869.8 | -2.2477 |
52869.8 | -2.22198 |
52870.0 | -1.95623 |
52871.3 | -3.1621 |
52871.8 | -2.42201 |
52872.0 | -3.70788 |
52872.2 | -1.82194 |
52873.7 | -2.73634 |
52875.3 | -3.50215 |
52875.7 | -2.8135 |
52875.8 | -2.76206 |
53007.5 | -3.22257 |
53008.4 | -1.79097 |
53009.2 | -0.487956 |
53010.7 | -3.07971 |
53011.2 | -2.37105 |
53013.4 | -3.7998 |
53013.7 | -3.38833 |
53013.8 | -3.28831 |
53014.3 | -2.49108 |
epoch [MJD] | Position Residual [mas] |
51946.5 | -1.84392 |
51946.5 | -1.81306 |
51946.8 | -1.31929 |
51947.0 | -1.01754 |
51980.5 | -1.72745 |
51980.5 | -1.69316 |
51981.2 | -2.54699 |
51981.3 | -2.46126 |
51981.3 | -2.39268 |
51981.4 | -2.31724 |
51981.4 | -2.25209 |
51981.7 | -1.84747 |
51982.0 | -1.27826 |
51982.0 | -1.25769 |
51982.1 | -1.23369 |
51982.1 | -1.1171 |
51982.2 | -1.00051 |
51982.2 | -0.973082 |
51982.3 | -0.856497 |
51983.2 | -1.45314 |
52179.3 | -2.25963 |
52179.3 | -2.22819 |
52179.3 | -2.17962 |
52179.4 | -2.10532 |
52179.4 | -2.08532 |
52179.7 | -1.54811 |
52179.9 | -1.20236 |
52179.9 | -2.84827 |
52180.0 | -2.81112 |
52180.0 | -2.77683 |
52180.0 | -2.40365 |
52180.0 | -2.72826 |
52180.5 | -1.9996 |
52180.5 | -1.95674 |
52180.5 | -1.87387 |
52180.6 | -1.80815 |
52180.6 | -1.77386 |
52180.8 | -1.43096 |
52181.1 | -1.01091 |
52181.8 | -1.48525 |
52308.3 | -3.63452 |
52308.4 | -3.57451 |
52308.5 | -3.29734 |
52308.6 | -3.16018 |
52308.6 | -3.1316 |
52308.7 | -2.98016 |
52308.8 | -2.92872 |
52308.8 | -2.85443 |
52309.0 | -2.62297 |
52309.1 | -2.48581 |
52309.1 | -2.43152 |
52309.1 | -2.41723 |
52309.2 | -2.18006 |
52309.3 | -2.11148 |
52309.3 | -2.07148 |
52309.7 | -1.4714 |
52309.8 | -2.90586 |
52309.9 | -1.12279 |
52310.0 | -0.977059 |
52310.1 | -0.854187 |
52310.3 | -0.559866 |
52311.3 | -0.51415 |
52497.0 | -3.48657 |
52497.3 | -3.02366 |
52498.0 | -3.63802 |
52498.3 | -3.20654 |
52498.4 | -3.01224 |
52498.5 | -2.78364 |
52499.5 | -2.92937 |
52499.8 | -2.37502 |
52503.6 | -3.07798 |
52503.7 | -2.86081 |
52503.7 | -2.82937 |
52504.0 | -2.45504 |
52505.0 | -2.44647 |
52505.4 | -3.54375 |
52506.3 | -2.06929 |
52508.5 | -3.62092 |
52509.5 | -1.955 |
52510.6 | -1.88357 |
52512.9 | -3.27803 |
52513.8 | -3.56664 |
52514.1 | -2.978 |
52514.3 | -2.64653 |
52535.5 | -3.98962 |
52535.7 | -3.68959 |
52535.7 | -3.65816 |
52536.2 | -2.90092 |
52537.5 | -2.4323 |
52539.7 | -4.01821 |
52539.8 | -3.82962 |
52540.2 | -3.16097 |
52540.7 | -2.3523 |
52542.4 | -3.06096 |
52543.7 | -2.58091 |
52544.0 | -2.20086 |
52642.1 | -2.79556 |
52642.9 | -3.17846 |
52642.9 | -3.1156 |
52643.8 | -1.664 |
52645.5 | -2.28979 |
52646.2 | -2.97845 |
52646.4 | -2.51268 |
52646.5 | -2.43553 |
52646.6 | -2.23836 |
52649.4 | -2.8413 |
52650.1 | -3.40137 |
52650.5 | -2.72129 |
52651.5 | -2.7613 |
52673.0 | -3.68148 |
52673.6 | -2.71851 |
52673.6 | -1.03259 |
52673.7 | -2.57278 |
52674.2 | -1.65267 |
52678.4 | -3.3929 |
52678.8 | -2.6871 |
52679.1 | -2.30419 |
52679.4 | -3.42719 |
52680.8 | -2.91285 |
52682.2 | -3.95012 |
52682.3 | -3.81868 |
52683.1 | -2.52423 |
52683.3 | -2.21277 |
52683.8 | -1.41553 |
52684.0 | -1.12121 |
52685.9 | -3.13575 |
52686.5 | -3.82726 |
52686.7 | -1.86702 |
52686.8 | -3.33006 |
52688.2 | -2.78428 |
52689.7 | -1.98133 |
52689.8 | -1.87846 |
52860.9 | -3.11348 |
52862.1 | -2.87346 |
52864.3 | -2.65916 |
52864.6 | -3.80215 |
52865.6 | -2.35055 |
52866.5 | -2.49057 |
52869.4 | -2.92206 |
52869.6 | -2.59917 |
52869.8 | -2.2477 |
52869.8 | -2.22198 |
52870.0 | -1.95623 |
52871.3 | -3.1621 |
52871.8 | -2.42201 |
52872.0 | -3.70788 |
52872.2 | -1.82194 |
52873.7 | -2.73634 |
52875.3 | -3.50215 |
52875.7 | -2.8135 |
52875.8 | -2.76206 |
53007.5 | -3.22257 |
53008.4 | -1.79097 |
53009.2 | -0.487956 |
53010.7 | -3.07971 |
53011.2 | -2.37105 |
53013.4 | -3.7998 |
53013.7 | -3.38833 |
53013.8 | -3.28831 |
53014.3 | -2.49108 |